Use elimination to solve:
3x-y=15
5x+y=9
this is what I got
3x-y=15
(+)5x+y=9
__________
8x+0=24
8x=24
8x/8= 24/8
x=3

3(2)-y=15
6-y=15
6-y-6=15-6
-y=9
-y/-1= 9/-1
y= -9
I do not know what I did wrong????????

Found 2 solutions by stanbon, wuwei96815:
Answer by stanbon(75887)

You correctly solved for "x", but you did
not correctly substitute that value into
the 1st original equation to solve for "y".
Cheers,
Stan H.

Answer by wuwei96815(245)

Elimination means to get rid of one of the variables so that you can solve for the other. With the first system of equations we have 3x-y=15 and 5x+y=9. Adding these two together we get 8x=24 (the -y and the +y added together eliminated the variable y). Now, it is a simple matter to solve for x.

8x=24
x = 24/8 = 3

We can now go back to the original equations and by substituting 3 for x solve for y. We can use either equation. Let's use the first one 3x-y=15.

3x-y=15
3(3)-y=15
9-y =15
-y=15-9=6
-y=6
y = -6

I think the mistake that you made was substituting (2) instead of (3) for the value of x. I hope this has helped you.